Please give the PROOFS of this equation!?!?
Statements
1. 3x - 7 = -4
2. 3x - 7 + 7 = -4 + 7
3. 3x + 0 = -4 + 7
4. 3x + 0 = 3
5. 3x = 3
6. (1/3) 3x = 3 (1/3)
7. (1/3) 3x = 1
8. 1x = 1
9. x = 1

Answer by drk(1908)

I think you are looking for the "reasons" after each "statement"
1. 3x - 7 = -4 REASON "given"
2. 3x - 7 + 7 = -4 + 7 REASON "addition property of equality"
3. 3x + 0 = -4 + 7 REASON "inverse property of addition"
4. 3x + 0 = 3 REASON "distributive property" - I know weird but true.
5. 3x = 3 REASON "identity property of addition"
6. (1/3) 3x = 3 (1/3) REASON "multiplicative inverse property" - you may have division in book.
7. (1/3) 3x = 1 REASON "any number, except 0, times its reciprocal is 1"
8. 1x = 1 REASON "any number, except 0, times its reciprocal is 1"
9. x = 1 REASON "identity property of addition"